# Sakib Matina, Alice Allena,b, Emily Shinklec, Aleksandra Pachalievad, Galen T. Cravena, Benjamin Nebgena, Justin Smithe, Richard Messerlya, Ying Wai Lib, Sergei Tretiaka,b,f, Kipton Barrosa,b, and Nicholas Lubbersc

### v1

@@ 3.2 Accuracy and MLIP Capacity
The force RMSE errors in panel (A) of Fig 3 indicate a meaningful improvement of the student models compared with the control models, while the energy errors in panel (B) are only minimally improved in the student models. We compute the RDFs of liquid aluminum at 1200 K using the LAVA 61 software, which is a wrapper for LAMMPS 5. The reported RDF errors shown in panel (C) for the student and control models are the total absolute errors with respect to the teacher RDFs. This out-of-sample MD test shows a similar trend as the MLIP training/test errors. Note that the energy and forces errors in Fig. 3(A) and (B) respectively are computed with respect to the ground truth training data, whereas the RDF errors in (C) are the differences against the teacher models’ MD simulations.
@@ 3.3 Learning Dynamics
Learning dynamics (also known as optimization dynamics) characterize how the training or out-of-sample validation error evolves as a function of the training epochs. Here, we analyze learning dynamics of the student and control MLIPs to understand the how the auxiliary targets affect the training. We see in Fig. 4 that the student models’ learning dynamics out perform the control models. loss curves of four different student and control models, which all share the same architecture but different initial weights. Consequently this is reflected in different intercepts, showing that student models train faster than the control models by a roughly constant factor.
@@ 3.4 Speed and Scalability
Our work shows that the teacher-student procedure has several benefits for large-scale MD. The light-weight student models require fewer floating point operations for force evaluations. Fig. 5 examines how the number of weights in the network affect the computational cost for large-scale MD in terms of both speed and memory in panels (A) and (B) respectively. The MD is performed using ASE in the NVE ensemble for 1 ps with a time step of 1 fs. Furthermore, we see that the smaller student models can fit more atoms on a single GPU (A6000 48 GB in Fig. 5). The ability to fit more atoms on a single GPU can improve the efficiency of large MD simulations, because inter-node latency can often dominate the computations 5.
@@ 3.5 Student MLIP Surpasses the Teacher
We apply the “Born Again” 46 teacher-student training to HIPNN. The student and teacher models have the same of trainable weights. We show that student models trained to the ground truth data and the teacher model’s auxiliary outputs and can surpass the teacher by using an loss scheduler.
Initially, we use the static loss function in Eq. 2 to find that student MLIPs achieve comparable errors to the teachers. However, then we utilize a loss scheduler to dynamically update the weights $w_{A},w_{F},w_{E}$ during the training as summarized in Table 1 in Appendix A.1. The final weights of the students’ loss function match that of that of the static weights of the teacher MLIP. The student models’ Energy RMSE of $0.37\pm 0.02\leavevmode\nobreak\ \mathrm{eV}$, and force RMSE of $0.083\pm 0.003\leavevmode\nobreak\ \mathrm{eV/$\mathrm{\SIUnitSymbolAngstrom}$}$ is lower than the teacher models errors, energy RMSE of $0.38\mathrm{eV}$, and force RMSE of $0.092\leavevmode\nobreak\ \mathrm{eV/$\mathrm{\SIUnitSymbolAngstrom}$}$. This shows that the BA teacher-student approach improves the force accuracy by about $10\%$, which shares the same model architecture and underlying training data.
@@ 4 Conclusions and Future work
We show that our teacher-student framework can reduce the inference costs of MLIPs and increase their accuracy without any additional QC data. Our teacher-student training framework improves the Pareto set of error-cost trade-offs for MLIPs. We show that the student MLIPs are Pareto dominant with respect to the control MLIPs. The student MLIPs, trained to the ground truth and teacher’s atomic energies, achieve higher accuracy at the same efficiency (speed and memory requirements) when compared to the control models that were only trained to ground truth. We use both training errors (energy and force RMSEs) as well as MD based metric (RDF errors) to quantify the accuracy of the MLIPs. We find that the force RMSE errors generally correlate with the MD based metrics, which are costlier to evaluate. Additionally, such MD based metrics require a ground truth, such as ab initio results or experimental data, which may not always be available. Efficiency of MLIPs is also multi-faceted. We use MD speeds and memory-per-atom as effective measures. While MD speed is easy to interpret, it strongly depends on the underlying hardware, MD codes as well as the system under study. The memory-per-atom measure is useful to optimize for large MD simulations, where inter-node communications may dominate. This teacher-student training framework can be readily applied to other MLIPs that decompose the configurational energy into a sum of local contributions.
We used the atomic energies as the knowledge for the teacher-student-training. In the language of message passing graph neural networks, the atom energy is a node level scalar 49. We intend to explore node level vectors, such as forces, and edge level properties as knowledge for the teacher-student framework to train across different architectures 43. Additionally, network weights for the interactions layers may be utilized as knowledge for the student. We will also explore using an ensemble of teachers for future work, where the students may be able to leverage the uncertainty associated with the teacher MLIPs’ knowledge.
Graph neural networks can be used to predict molecular properties beyond just energy and forces such as charges 62, 29, dipoles 28, and other properties 11. We will explore cross-modal teacher-student training frameworks to combine teacher models with different specialized tasks to train generalist student MLIPs. This will be an important step towards foundation models for chemistry with broad applicability.
@@ A.1 Hyper-Parameters
We use HIPNN models with $1$ interaction layer. The teacher models use $4$ atom layers (feed-forward layers) with a width of $128$. The student (and control) models have between $1$ and $4$ atom layers with width between $12$ and $128$. All models have maximum tensor sensitivity order set at $l_{\mathrm{max}}=1$. For the sensitivity functions which parameterize the interaction layer, radial basis functions are used with a soft-min cutoff of $1.5\leavevmode\nobreak\ $\mathrm{\SIUnitSymbolAngstrom}$$, the soft maximum cutoff of $7.0\leavevmode\nobreak\ $\mathrm{\SIUnitSymbolAngstrom}$$, and hard maximum cutoff of $7.5\leavevmode\nobreak\ $\mathrm{\SIUnitSymbolAngstrom}$$. The teacher model uses $40$ basis functions. For the student (and control) MLIPs, we use between $8$ and $40$ sensitivity functions. The soft-min cut-off corresponds to the inner cut-off at very short distances. The hard maximum cut-off corresponds to the long distance cut-off. The soft maximum cutoff is set to a value smaller than the hard-dist cutoff to ensure a smooth truncation of the sensitivity functions. Note that the we are using the naming conventions for the hyper-parameters in the HIPNN GitHub Repository 63, which differs slightly from the original HIPNN publication 22.
@@
The HIPNN 22, 14 MLIP is implemented in an open-source PyTorch-based software package called hippynn which is available for download 63.
We use an aluminum dataset published in Ref. 23 which is available for download 59.
@@ Acknowledgements
This work was supported by the United States Department of Energy (US DOE), Office of Science, Basic Energy Sciences, Chemical Sciences, Geosciences, and Biosciences Division under Triad National Security, LLC (‘Triad’) contract grant no. 89233218CNA000001 (FWP: LANLE3F2). We acknowledge support from the Los Alamos National Laboratory (LANL) Directed Research and Development funds (LDRD). This research was performed in part at the Center for Nonlinear Studies (CNLS) at LANL. This research used resources provided by the Darwin testbed at LANL which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program.
@@
  • Giannozzi et al. 2017 P. Giannozzi, O. Andreussi, T. Brumme, O. Bunau, M. B. Nardelli, M. Calandra, R. Car, C. Cavazzoni, D. Ceresoli, M. Cococcioni, N. Colonna, I. Carnimeo, A. D. Corso, S. de Gironcoli, P. Delugas, R. A. DiStasio, A. Ferretti, A. Floris, G. Fratesi, G. Fugallo, R. Gebauer, U. Gerstmann, F. Giustino, T. Gorni, J. Jia, M. Kawamura, H.-Y. Ko, A. Kokalj, E. Küçükbenli, M. Lazzeri, M. Marsili, N. Marzari, F. Mauri, N. L. Nguyen, H.-V. Nguyen, A. O. de-la Roza, L. Paulatto, S. Poncé, D. Rocca, R. Sabatini, B. Santra, M. Schlipf, A. P. Seitsonen, A. Smogunov, I. Timrov, T. Thonhauser, P. Umari, N. Vast, X. Wu and S. Baroni, Journal of physics: Condensed matter, 2017, 29, 465901.
  • 59 ANI-AL data and model GitHub repository. https://github.com/atomistic-ml/ani-al.
@@
  • 63 The hippynn package - a modular library for atomistic machine learning with PyTorch, available online: https://github.com/lanl/hippynn.

### v2

@@ 3.2 Accuracy and MLIP Capacity
The force RMSE errors in panel (A) of Fig 3 indicate a meaningful improvement of the student models compared with the control models, while the energy errors in panel (B) are only minimally improved in the student models. We compute the RDFs of liquid aluminum at 1200 K using the LAVA 65 software, which is a wrapper for LAMMPS 5. The reported RDF errors shown in panel (C) for the student and control models are the total absolute errors with respect to the teacher RDFs. Note that the energy and forces errors in Fig. 3(A) and (B) respectively are computed with respect to the ground truth training data, whereas the RDF errors in (C) are the differences against the teacher models’ MD simulations. Figure 3 reveals that the energy RMSE is quite similar between the student and control models, except for very small model sizes. We attribute the improvements in the force accuracy to the fact that the atomic energies are local properties predicted for each atom, similar to atomic forces. Furthermore, the force errors strongly correlate with the RDF errors. Note that the RDF simulations constitute a difficult extensibility test, because the simulations utilize large periodic boxes that are two orders of magnitude larger than the configurations found in the training data, with each simulation containing 18634 atoms.
@@ 3.3 Learning Dynamics
Learning dynamics (also known as optimization dynamics) characterize how the training or out-of-sample validation error evolves as a function of the training epochs. We analyze learning dynamics of the student and control MLIPs to understand the how the auxiliary targets affect the training. We see in Fig. 4 that the student models’ learning dynamics out perform the control models. Loss curves of four different student and control models, which all share the same architecture. Here, we use four different seed values for the random initialization of the sets of four control and four students respective. By employing the same seed for a pair of student and control MLIPs, we ensure that each student MLIP has the same initial weights as its corresponding control MLIP. Note that the learning dynamics curves begin after one epoch of training has been done. The apparent improved initialization of the student models are actually due to the improved learning by the student HIPNN models due to the atomic energy targets from the teacher. We also note that the auxiliary targets have a regularization effect, stabilizing the optimization, because the learning curves of the student models exhibit less variance in the later stages of training.
@@ 3.4 Speed and Scalability
Our work shows that the teacher-student procedure lowers the resource required for large-scale MD at a target accuracy. The teacher-student method improves the accuracy of smaller models. The light-weight student models require fewer floating point operations for force evaluations. As a direct consequence, we can now run large scale MD at a desired accuracy while using fewer computational resources. Fig. 5 examines how the number of weights in the network affect the computational cost for large-scale MD in terms of both speed and memory in panels (A) and (B) respectively. The MD is performed using ASE in the NVE ensemble for 1 ps with a time step of 1 fs. Furthermore, we see that the smaller student models can fit more atoms on a single GPU (A6000 48 GB in Fig. 5). The ability to fit more atoms on a single GPU can improve the efficiency of large MD simulations, because inter-node latency can often dominate the computations 5.
@@ 3.5 Student MLIP Surpasses the Teacher
We apply the “Born Again” 47 teacher-student training to HIPNN. The ‘Born Again’ method was introduced in the Ref 47, and it is a variant of the teacher-student training where both models have exact same architecture and number of trainable weights. The central aim of the ‘Born Again’ method is to train a student model that surpasses the teacher’s accuracy. This is in contrast to knowledge distillation, where the aim is to improve the accuracy of smaller student models that have fewer trainable weights than the teacher models. We show that student models trained to the ground truth data and the teacher model’s auxiliary outputs and can surpass the teacher by using an loss scheduler.
Initially, we use the static loss function in Eq. 2 to find that student MLIPs achieve comparable errors to the teachers. However, then we utilize a loss scheduler to dynamically update the weights $w_{A},w_{F},w_{E}$ during the training as summarized in Table 1 in Appendix A.1. The weights in the loss function were chosen to favor the teacher’s knowledge in the early stages of the training and emphasizing the QM data in the later stages, so that the student can surpass the accuracy on the QM. The values of the weight schedule were determined through manual hyper parameter tuning. The final weights of the students’ loss function match that of that of the static weights of the teacher MLIP. The student models’ Energy RMSE of $0.37\pm 0.02\leavevmode\nobreak\ \mathrm{eV}$, and force RMSE of $0.083\pm 0.003\leavevmode\nobreak\ \mathrm{eV/$\mathrm{\SIUnitSymbolAngstrom}$}$ is lower than the teacher models errors, energy RMSE of $0.38\mathrm{eV}$, and force RMSE of $0.092\leavevmode\nobreak\ \mathrm{eV/$\mathrm{\SIUnitSymbolAngstrom}$}$. This shows that the BA teacher-student approach improves the force accuracy by about $10\%$, which shares the same model architecture and underlying training data.
@@ 4 Conclusions and Future work
We introduce a teacher-student framework that can be readily applied to many MLIPs that decompose the configurational energy into a sum of local contributions. The teacher-student training framework improves the Pareto set of error-cost trade-offs for MLIPs, yielding models that are computationally cheaper for the same accuracy, or more accurate for the same computational cost, and requires no additional ground-truth data.
In a practical setting, we showed that the student MLIPs are Pareto dominant with respect to the control MLIPs. The student MLIPs, trained to the ground truth and teacher’s atomic energies, achieve higher accuracy at the same efficiency (speed and memory requirements) when compared to the control models that were only trained to ground truth. We use both training errors (energy and force RMSEs) as well as MD based metric (RDF errors) to quantify the accuracy of the MLIPs. We find that the force RMSE errors generally correlate with the MD based metrics, which are costlier to evaluate. MD based metrics also demonstrate the extensibility of the MLIP to configurations much larger than what is typically included in the training dataset. However, MD based metrics require a ground truth, such as ab initio results or experimental data, which may not always be available. Efficiency of MLIPs is also multi-faceted. We use MD speeds and memory-per-atom as effective measures. While MD speed is easy to interpret, it strongly depends on the underlying hardware, MD codes, and the system under study. The memory-per-atom measure is useful to optimize for large MD simulations, where inter-node communications may dominate. While the research to improve the accuracy of MLIPs has focused on the development of more expressive architectures 50, and the generation of larger datasets 66, we show that the ‘born-again’ inspired teacher-student training can be used to train more accurate models with existing datasets and architectures. Thus, innovations to the training protocol can allow us to extract better models from existing datasets.
We used the atomic energies as the knowledge for the teacher-student-training. In the language of message passing graph neural networks, the atom energy is a node level scalar 50. In future works, one can explore node level vectors, such as forces 56, 57, and edge level properties as knowledge for the teacher-student framework to train across different architectures 43, 44. Additionally, network weights for the interactions layers may be utilized as knowledge for the student. One may also explore using an ensemble of teachers for future work, where the students may be able to leverage the uncertainty associated with the teacher MLIPs’ knowledge.
Graph neural networks can be used to predict molecular properties beyond just energy and forces such as charges 67, 29, dipoles 28, and other properties 11. There is potential to explore cross-modal teacher-student training frameworks to combine teacher models with different specialized tasks to train generalist student MLIPs. This will be an important step towards foundation models for chemistry with broad applicability.
@@ A.1 Hyper-Parameters
We use HIPNN models with $1$ interaction layer. The teacher models use $4$ atom layers (feed-forward layers) with a width of $128$. The student (and control) models have between $1$ and $4$ atom layers with width between $12$ and $128$. All models have maximum tensor sensitivity order set at $l_{\mathrm{max}}=1$. For the sensitivity functions which parameterize the interaction layer, radial basis functions are used with a soft-min cutoff of $1.5\leavevmode\nobreak\ $\mathrm{\SIUnitSymbolAngstrom}$$, the soft maximum cutoff of $7.0\leavevmode\nobreak\ $\mathrm{\SIUnitSymbolAngstrom}$$, and hard maximum cutoff of $7.5\leavevmode\nobreak\ $\mathrm{\SIUnitSymbolAngstrom}$$. The teacher model uses $40$ basis functions. For the student (and control) MLIPs, we use between $8$ and $40$ sensitivity functions. The soft-min cut-off corresponds to the inner cut-off at very short distances. The hard maximum cut-off corresponds to the long distance cut-off. The soft maximum cutoff is set to a value smaller than the hard-dist cutoff to ensure a smooth truncation of the sensitivity functions. Note that the we are using the naming conventions for the hyper-parameters in the HIPNN GitHub Repository 68, which differs slightly from the original HIPNN publication 22.
@@
The HIPNN 22, 14 MLIP is implemented in an open-source PyTorch-based software package called hippynn which is available for download 68. We will upload the scripts for workflow and trained models to the examples folder. We use an aluminum dataset published in Ref. 23 and the dataset is available for download 63.
Appendix B Author contributions
Conceptualization (SM, NL); Methodology (SM, AEAA, ES, AP, GTC, BN, JSS, RM, YWL, ST, KB, NL); Formal analysis (SM); Writing - original draft (SM); Writing - review and editing (SM, AEAA, ES, AP, GTC, BN, JSS, RM, YWL, ST, KB, NL); Supervision (GTC, NL).
@@ Acknowledgements
This work was supported by the United States Department of Energy (US DOE), Office of Science, Basic Energy Sciences, Chemical Sciences, Geosciences, and Biosciences Division under Triad National Security, LLC (‘Triad’) contract grant no. 89233218CNA000001 (FWP: LANLE3F2, LANLE8AN). We acknowledge support from the Los Alamos National Laboratory (LANL) Directed Research and Development funds (LDRD). This research was performed in part at the Center for Nonlinear Studies (CNLS) at LANL. This research used resources provided by the Los Alamos National Laboratory Institutional Computing Program and Darwin testbed at LANL which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program.
@@
  • Amin et al. 2025 I. Amin, S. Raja and A. Krishnapriyan, arXiv preprint arXiv:2501.09009, 2025.
@@
  • Gong et al. 2025 S. Gong, Y. Zhang, Z. Mu, Z. Pu, H. Wang, X. Han, Z. Yu, M. Chen, T. Zheng, Z. Wang et al., Nature Machine Intelligence, 2025, 1–10.
  • Matin et al. 2025 S. Matin, E. Shinkle, Y. Pimonova, G. T. Craven, A. Pachalieva, Y. W. Li, K. Barros and N. Lubbers, Ensemble Knowledge Distillation for Machine Learning Interatomic Potentials, 2025, https://arxiv.org/abs/2503.14293.
@@
  • F. dos Santos et al. 2025 L. G. F. dos Santos, B. T. Nebgen, A. E. A. Allen, B. W. Hamilton, S. Matin, J. S. Smith and R. A. Messerly, Journal of Chemical Information and Modeling, 2025.
@@
  • Jung 2024 G. S. Jung, Journal of Chemical Information and Modeling, 2024.
  • 63 ANI-AL data and model GitHub repository. https://github.com/atomistic-ml/ani-al.
@@
  • Kaplan et al. 2025 A. D. Kaplan, R. Liu, J. Qi, T. W. Ko, B. Deng, J. Riebesell, G. Ceder, K. A. Persson and S. P. Ong, A Foundational Potential Energy Surface Dataset for Materials, 2025, https://arxiv.org/abs/2503.04070.
@@
  • 68 The hippynn package - a modular library for atomistic machine learning with PyTorch, available online: https://github.com/lanl/hippynn.